# Enhancing Language Model Reasoning via Weighted Reasoning in Self-Consistency

### v1

@@ Abstract
While large language models (LLMs) have rapidly improved their performance on a broad number of tasks, they still often fall short on reasoning tasks. As LLMs become more integrated in diverse real-world tasks, advancing their reasoning capabilities is crucial to their effectiveness in nuanced, complex problems. Wang et al. [27]’s self-consistency framework reveals that sampling multiple rationales before taking a majority vote reliably improves model performance across various closed-answer reasoning tasks. Standard methods based on this framework aggregate the final decisions of these rationales but fail to utilize the detailed step-by-step reasoning paths applied by these paths. Our work enhances this approach by incorporating and analyzing both the reasoning paths of these rationales in addition to their final decisions before taking a majority vote. These methods not only improve the reliability of reasoning paths but also cause more robust performance on complex reasoning tasks.
@@ 2 Datasets
We evaluate the models on arithmetic and commonsense reasoning using three datasets: AQuA-RAT, SVAMP, and StrategyQA. AQuA-RAT assesses models’ ability to solve arithmetic problems involving basic calculations, numerical relationships, and multi-step reasoning [14]. SVAMP challenges models with math problems focused on algebraic manipulations and symbolic reasoning [21]. StrategyQA tests models on answering complex, open-domain questions that require strategic thinking rather than simple factual knowledge [8].
@@ 3 Language Models
Detailed information on the configurations used for our models can be found in Appendix E. We use chain-of-thought prompting for all of our experiments. The prompts can be found in Appendix G.
@@ 3.2 Featurizers
All of our featurizers are based on the BERT (Bidirectional Encoder Representations from Transformers) model architecture [6], with various fine-tuned versions used to generate embedding vectors tailored to specific datasets. RoBERTa is employed for the StrategyQA dataset, which requires reading comprehension and contextual reasoning, benefiting from RoBERTa’s robustness in general language processing tasks [16]. Additionally, we use SciBERT for the AQuA-RAT and SVAMP datasets, which focus on mathematical reasoning, as its specialization in scientific texts makes it well-suited to handle the language present in these datasets [2].
@@ 5.1 Semantic consistency results
CPW improved self-consistency by 3.14% on AQuA-RAT and 0.97% on SVAMP but decreased performance by -1.63% on StrategyQA, likely due to its limited reasoning paths. This effect was seen across self-consistency, where improvements were smaller compared to other datasets. A detailed discussion of these suboptimal cases is in Appendix A.3.
@@ 5.2 Outlier detection results
The results from our analysis of various outlier detection methods—isolation forest, k-nearest neighbor, one-class support vector machines (SVM)—demonstrate their effectiveness in refining the quality of model output. The observed increases in accuracy across these methods remain consistent towards reduced sample sizes as well, suggesting that the effectiveness of anomaly detection techniques are not solely dependent on sample size. Obtained results exhibited slight deviations between the different configurations. A review across different sets of configurations and parameters can be found under Appendix J.1 to J.3.
@@
Anomaly detection serves a diagnostic tool to reduce the sample space particularly in noisy scenarios. Additionally methods generally improved performance over the SC baseline. AQuA-RAT saw an average improvement of 8.27%, while StrategyQA had a modest 1.33% gain. However, SVAMP showed a slight decline of 1.99%. These results highlight variability across datasets, with isolation forest and one-class SVM performing better on certain datasets.
@@ 6 Discussion
It is worth noting that our system uses embedding vectors to filter responses based on general reasoning accuracy, prioritizing broad similarity over subtle variations, as the benefit of choosing the numerical majority vote from self-consistency to yield correct answers still applies, especially in the limited rationale space. An additional analysis can be found in Appendix A.1.
Diverse responses are not necessarily undesirable and can lead to elevated results as shown in Appendix C.1. Against the natural feel, employed methods do not discriminate against diverse reasoning. Lowering the temperature will make multiple responses more diverse and, therefore, broaden the distribution. This will not affect performance when using CPW or outlier detection, since outputs farther from the mean are not outliers but sensible parts of a wider distribution. Consequently, the weighting process will remain consistent, as all values will proportionally receive lower weights.
@@ 7 Conclusion
Our investigation into weighting and anomaly detection methods shows that cosine similarity outperforms CPW in improving model accuracy, particularly for models like Llama 2 7B and GPT 3.5 on datasets such as StrategyQA. CPW was effective for AQuA-RAT and SVAMP, leading to accuracy increases, but less so for StrategyQA. Our system prioritizes general reasoning accuracy using embedding vectors, with numerical majority voting from self-consistency remaining a key factor in achieving correct answers, especially within limited rationale spaces.
@@ 10 Ethical Considerations & Risks
Language models may produce factually incorrect information or biased outputs based on user prompts. The BERT-based featurizers, trained only on English corpora, may yield inconsistent or biased results when used with other languages. Mistral 7B, Llama 2 7B, and Llama 3 8B lack built-in content moderation, requiring external safeguards to avoid harmful content. While GPT-4o and GPT-4o mini have stronger moderation, biases may still surface.
We encourage responsible use of these models and the implementation of monitoring tools to mitigate potential risks.
Appendix A Additional Studies
A.1 Effects of symbolic logic and embeddings
@@
A.2 Embedding quality analysis
It is important to distinguish that the employed system focuses on identifying consensual responses and broader similarity in the representational space of embeddings, rather then subtle nuances. A clear analysis of our embeddings in connection to symbolic logic and subtle details can be found in Appendix A.1.
@@ A.2 Embedding quality analysis
Further analysis of both the embedding distribution as well as our dimensionality reduction can be found in Appendix L.
A.3 Self-consistency failure scenarios
@@
  • Overly similar generations: Generations that provide overly similar reasoning will likely be categorized in a similiar position in the embedding space, which will lead to our semantic methods, not being able to discern between elements.
  • Small subtleties in generations: As described in Section A.1 & Section 6 small subtleties aren’t captured directly by our model, making it less capable in tasks like Symbolic Reasoning or state tracking.
A.4 Comparison and effects
Meta-reasoning over multiple chains-of-thought While meta-reasoning has proven effective on tasks that have qualitative evident information, its ability to stay consistent between arithmetic operations and its subsequent reasoning path witnesses the same limitations as baseline self-consistency and chain-of-thought [29].
Appendix B Sample analysis
In the evaluation of Aquarat, some results exhibited noise. Particularly smaller models failed to consistently follow the few-shot chain-of-thought examples occasionally. This led to instances where outputs could not be parsed for final analysis. To ensure reproducibility, we employed the parsing extraction approach from baseline self-consistency. Furthermore, some models showed degeneration after generating the initial response, highlighting the need of development for a custom extraction function to ensure accurate semantic interpretation, particularly when utilizing functions that include embeddings.
Appendix C Performance variation
Across different findings, we see a variation in performance with a general upward trend. As shown in Section 3 and discussed in Appendix A.1, sequence length seems to affect model performance positively. Smaller sequences tend to contain less information, as seen with LLAMA 7B’s position in the table.
C.1 Sampling from multiple temperatures
@@
It is to note that higher temperature showed a degree of randomness that can lead to higher degeneration. However this limiting factor can be mitigated when applied with inverse temperature weighting and improve performance of up to 2.5%. The effect of different temperature sets can be found in Appendix K.
C.2 Finetuned featurizers
@@
C.3 Secondary semantic evaluation methods
@@
This method implies that the predictions associated with the majority cluster are the ones for which the model exhibits the greatest overall confidence. A detailed assessment of the found results can be accessed in Appendix I.1.
Appendix D N-Gram Rationale Comparison
D.1 Rouge-N as a performance measure
@@ D.2 N-Gram weighting
Model AQUARAT SVAMP
@@
Appendix E Model configurations
Configurations may deviate slightly on GPT 3.5 due to usage via the public API.
E.1 max-new-tokens
We used a default of 150 max-new-tokens across all models on SVAMP, due to the complexity and length of sequences on AQuA-RAT we chose 200 max-new-tokens. Due to the length of code tasks we set the max generation of new tokens to 400 on humaneval.
Appendix F Abstract consistency
F.1 Temperature sets
@@
F.2 Weighing abstract consistency
@@ Appendix G Prompting
G.1 Used k-shot prompts
@@
A: Leah had 32 chocolates and Leah’s sister had 42. That means there were originally 32 + 42 = 74 chocolates. 35 have been eaten. So in total they still have 74 - 35 = 39 chocolates. The answer is 39. Q: Jason had 20 lollipops. He gave Denny some lollipops. Now Jason has 12 lollipops. How many lollipops did Jason give to Denny?
@@
A: Michael initially had 58 balls. He lost 23 on Tuesday, so after that he has 58 - 23 = 35 balls. On Wednesday he lost 2 more so now he has 35 - 2 = 33 balls. The answer is 33.
@@
A: She bought 5 bagels for \$3 each. This means she spent 5 * \$3 = \$15 on the bagels. She had \$23 in beginning, so now she has \$23 - \$15 = \$8. The answer is 8
Proposed 4-shot on AQuA(-rat):
@@
A: If 10 is added to each number, then the mean of the numbers also increases by 10. So the new mean would be
50. The answer is (a).
@@
Q: A person is traveling at 20 km/hr and reached his destiny in 2.5 hr then find the distance?
@@
Our generation on humaneval were conducted 0-shot using just the raw prompt given by the dataset.
Appendix H Datasets
We use the configuration splits for testing as suggested by default. We employ the train and test split with 1000 samples for a less noisy assessment on SVAMP. For AQuA-RAT, our test includes the full set of 254 examples. With StrategyQA we use the full test split of 687 samples
Appendix I K-means Clustering
@@ I.1.1 Silouhette score
I.1.2 Average correct datapoint proportion
@@
Table 11: Proportion of correct responses in the majority cluster compared to total true responses.
Model SVAMP AQuA-RAT
Llama 2 68.8 56.6
Mistral 66.2 46.2
GPT 3.5 69.4 55.5
@@
I.1.3 Cluster density comparison
@@
SVAMP GPT 4.428 4.513
AQuA-RAT LLAMA 0.838 0.670
@@
Appendix J Outlier detection across different parameters
J.1 k-nearest neighbor results
In the k-nearest neighbor (KNN) algorithm, parameters such as the number of neighbors (n_neighbors), the distance metric (metric), and the algorithm used for computing nearest neighbors were varied. The best-performing configuration in terms of accuracy was found with n_neighbors set to 5, using the euclidean metric using the ball_tree algorithm and a threshold of 90% that concluded to an averaged accuracy of 56.18% with all Models and Datasets.
J.2 Isolation forest results
J.3 support vector machines results
Appendix K Abstract consistency on different temperature sets
@@
Appendix L Dimensionality reduction
@@
Appendix M Appendix / supplemental material
Optionally include supplemental material (complete proofs, additional experiments and plots) in appendix. All such materials SHOULD be included in the main submission.
@@
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.

### v2

# Semantic Self-Consistency: Enhancing Language Model Reasoning via Semantic Weighting

@@ Abstract
While large language models (LLMs) have rapidly improved their performance on a broad number of tasks, they still often fall short on reasoning tasks. As LLMs become more integrated in diverse real-world tasks, advancing their reasoning capabilities is crucial to their effectiveness in nuanced, complex problems. Wang et al. [33]’s self-consistency framework reveals that sampling multiple rationales before taking a majority vote reliably improves model performance across various closed-answer reasoning tasks. Standard methods based on this framework aggregate the final decisions of these rationales but fail to utilize the semantic information detailed in the step-by-step reasoning paths. Our work introduces semantic self-consistency, enhancing this approach by incorporating and analyzing both the reasoning paths of these rationales in addition to their final decisions before taking a majority vote. These methods not only improve the reliability of reasoning paths but also cause more robust performance on complex reasoning tasks.
@@ 2 Datasets
We evaluate the models on arithmetic and commonsense reasoning using three datasets: AQuA-RAT, SVAMP, and StrategyQA. AQuA-RAT assesses models’ ability to solve arithmetic problems involving basic calculations, numerical relationships, and multi-step reasoning [17]. SVAMP challenges models with math problems focused on algebraic manipulations and symbolic reasoning [25]. StrategyQA tests models on answering complex, open-domain questions that require strategic thinking rather than simple factual knowledge [9]. For specific information please refer to Appendix L.
@@ 3 Language Models
Detailed information on the configurations used for our models can be found in Appendix I.3. Additionally important hyperparameters for different methods are discussed in Appendix I We use chain-of-thought prompting for all of our experiments. The prompts can be found in Appendix K.
@@ 3.2 Featurizers
All of our featurizers are based on the BERT (Bidirectional Encoder Representations from Transformers) model architecture [7], with various fine-tuned versions used to generate embedding vectors tailored to specific datasets.RoBERTa 125M is employed for the StrategyQA dataset, which requires reading comprehension and contextual reasoning, benefiting from RoBERTa’s robustness in general language processing tasks [19]. Additionally, we use SciBERT 110M for the AQuA-RAT and SVAMP datasets, which focus on mathematical reasoning, as its specialization in scientific texts makes it well-suited to handle the language present in these datasets [2].
@@ 5.1 Semantic consistency results
CPW improved self-consistency by 3.14% on AQuA-RAT and 0.97% on SVAMP but decreased performance by -1.63% on StrategyQA, likely due to its limited reasoning paths. This effect was seen across self-consistency, where improvements were smaller compared to other datasets. A detailed discussion of these suboptimal cases is in Appendix D.
@@ 5.2 Outlier detection results
The results from our analysis of various outlier detection methods isolation forest, k-nearest neighbor, one-class support vector machines (SVM) demonstrate their effectiveness in refining the quality of model output. The observed increases in accuracy across these methods remain consistent towards reduced sample sizes as well, suggesting that the effectiveness of anomaly detection techniques are not solely dependent on sample size. Obtained results exhibited slight deviations between the different configurations. A review across different sets of configurations and parameters can be found under Appendix I.2.1 to I.2.3.
@@
The found results highlight variability across datasets, with isolation forest and one-class SVM performing better on certain datasets.
@@ 6 Discussion
It is worth noting that our system uses embedding vectors to filter responses based on general reasoning accuracy, prioritizing broad similarity over subtle variations, as the benefit of choosing the numerical majority vote from self-consistency to yield correct answers still applies, especially in the limited rationale space. An additional analysis can be found in Appendix B.
Diverse responses are not necessarily undesirable and can lead to elevated results as shown in Appendix G.1. Against the natural feel, employed methods do not discriminate against diverse reasoning. Lowering the temperature will make multiple responses more diverse and, therefore, broaden the distribution. This will not affect performance when using CPW or outlier detection, since outputs farther from the mean are not outliers but sensible parts of a wider distribution. Consequently, the weighting process will remain consistent, as all values will proportionally receive lower weights.
@@ 7 Conclusion
Our investigation into weighting and anomaly detection methods shows that cosine similarity outperforms CPW in improving model accuracy, particularly for models like Llama 2 7B and GPT 3.5 on datasets such as StrategyQA. CPW was effective for AQuA-RAT and SVAMP, leading to accuracy increases, but less so for StrategyQA. Our system prioritizes general reasoning accuracy using embedding vectors, with numerical majority voting from self-consistency remaining a key factor in achieving correct answers, especially within limited rationale spaces. Please note that the recommended methods should be employed with carefully tested hyperparameters, as their effectiveness may vary with subtle implementation nuances.
8 Related Work
Reasoning is an ubiquitous issue across many domains. [6]. One significant advancement in the area has been the development of the chain-of-thought prompting [34, 27] and self-consistency [33], which we extend for our Method. Self improvement of Language Models after generation is a well-known method for improving accuracy [12]. This concept has often been adapted by other weighting methods during pre-training to improve overall accuracy [31, 20], using different methods to shift the distribution [14].
@@ 11 Ethical Considerations & Risks
Language models may produce factually incorrect or biased outputs based on user prompts. The BERT-based featurizers, trained on English corpora, may yield inconsistent results in other languages. Mistral 7B, Llama 2 7B, and Llama 3 8B lack built-in content moderation, needing external safeguards against harmful content. While GPT-4o and GPT-4o mini have stronger moderation, biases may still emerge.
Further risks include that embedding and clustering methods may introduce subtle biases by emphasizing specific response types over others. Additionally variations in model temperature and sampling can add unintended randomness. Controlled sampling and inverse temperature weighting help but require careful tuning.
We recommend using monitoring tools and responsible model deployment, particularly in high-stakes applications.
12 Acknowledgements
We thank Celine Lee and Andy Chung for their helpful feedback on the methods and assistance in improving the clarity of our work. We also thank the anonymous reviewers of the conference for their insightful comments and suggestions, which helped improve the quality of this paper.
Appendix A Performance variation
Across different findings, we see a variation in performance with a general upward trend. As shown in Section 3 and discussed in Appendix B, sequence length seems to affect model performance positively. Smaller sequences tend to contain to be less similar in terms of informational density compared to all other sequences.
Appendix B Effects of symbolic logic and embeddings
@@
Appendix C Embedding quality analysis
It is important to distinguish that the employed system focuses on identifying consensual responses and broader similarity in the representational space of embeddings, rather then subtle nuances. A clear analysis of our embeddings in connection to symbolic logic and subtle details can be found in Appendix B.
@@ Appendix C Embedding quality analysis
Further analysis of both the embedding distribution as well as our dimensionality reduction can be found in Appendix O.
Appendix D Self-consistency failure scenarios
@@
  • Overly similar generations: Generations that provide overly similar reasoning will likely be categorized in a similar position in the embedding space, which will lead to our semantic methods, not being able to discern between elements.
  • Small subtleties in generations: As described in Section B & Section 6 small subtleties aren’t captured directly by our model, making it less capable in tasks like Symbolic Reasoning or state tracking.
Appendix E Comparison to related Methods
E.1 Meta-reasoning over multiple chains-of-thought
While meta-reasoning has proven effective on tasks that have qualitative evident information, its ability to stay consistent between arithmetic operations and its subsequent reasoning path witnesses the same limitations as baseline self-consistency and chain-of-thought [35].
E.2 Importance Weighting with self-improvement
Unlike previously established self-improvement and Importance weighting methods as proposed by Jiang et al. [14]. Our methods weighs results directly after generation in a separate weighting/filtering step. While results showed some frailty if not tuned with fitting parameters we spare computational efforts by not requiring an addition pre-training step. Pre-trained self-improvement Models could be used together with our introduced weighting method, to test performance and facilitate accuracy even further
Appendix F Sample analysis
In the evaluation of AQuA-RAT, some results exhibited noise. Particularly smaller models failed to consistently follow the few-shot chain-of-thought examples occasionally. This led to instances where outputs could not be parsed for final analysis. To ensure reproducibility, we employed the parsing extraction approach from baseline self-consistency. Furthermore, some models showed degeneration after generating the initial response, highlighting the need of development for a custom extraction function to ensure accurate semantic interpretation, particularly when utilizing functions that include embeddings.
Appendix G Efficiency Comparison
Other than self-consistency our methods require additional computation, other than the initial generation to compute its results.
  • Embeddings: The computational cost is moderate, as the BERT model utilized is of a manageable size, keeping resource usage at a reasonable level.
  • Centroid Proximity Weighting: This method is computationally inexpensive, as it relies solely on mathematical operations without requiring extensive resources.
  • Semantic Consensus Weighting: Similarly, this technique is computationally efficient due to its reliance on lightweight mathematical computations.
  • Outlier Detection: All three outlier detection methods employed are computationally low-cost, ensuring minimal impact on overall performance.
Compared to baseline self-consistency, the performance loss is minimal on modern GPUs, with the most computational effort still lying on the initial generation. Additionally, unlike other self-improvement methods, our techniques don’t require an extra pre-training phase and can be applied directly. As advancements in computational resources continue and smaller models grow increasingly capable, we expect this concern to become even less significant.
G.1 Sampling from multiple temperatures
@@
It is to note that higher temperature showed a degree of randomness that can lead to higher degeneration. However this limiting factor can be mitigated when applied with inverse temperature weighting and improve performance of up to 2.5%. The effect of different temperature sets can be found in Appendix N.
G.2 Finetuned featurizers
@@
G.3 Secondary semantic evaluation methods
@@
This method implies that the predictions associated with the majority cluster are the ones for which the model exhibits the greatest overall confidence. A detailed assessment of the found results can be accessed in Appendix M.1.
Appendix H N-Gram Rationale Comparison
H.1 Rouge-N as a performance measure
@@ H.2 N-Gram weighting
Model AQuA-RAT SVAMP
@@
Appendix I Configuration & Parameters
I.1 Varying Response Count
Our analysis indicates that maintaining a minimum of 7-10 responses is crucial to achieving consistent performance comparable to the baseline. When $k$ is set to lower values, the performance gains diminish, sometimes leading to completely random results with low accuracy. We expect that increasing the number of responses could enhance the effectiveness of our methods, leveraging the additional context and range of responses available to each weighting mechanism for improved accuracy. For all experiments we use $k$ = 10 unless stated otherwise.
I.2 Outlier detection Hyperparameters
I.2.1 k-nearest neighbor results
In the k-nearest neighbor (KNN) algorithm, parameters such as the number of neighbors (n_neighbors), the distance metric (metric), and the algorithm used for computing nearest neighbors were varied. The best-performing configuration in terms of accuracy was found with n_neighbors set to 5, using the euclidean metric using the ball_tree algorithm and a threshold of 90% [^3] To use KNN as an outlier detection system adapt the system for detecting outliers, by defining a threshold that retains part of the distribution based on the distance to nearest neighbors that concluded to an averaged accuracy of 56.18% with all Models and Datasets.
I.2.2 Isolation forest results
I.2.3 support vector machines results
I.3 Model configuration
Configurations may deviate slightly on GPT 3.5 & GPT4o-mini due to usage via the public API.
I.4 Token generation
We used a default of 250 max-new-tokens across all models on SVAMP, due to the complexity and length of sequences on AQuA-RAT we chose 400 max-new-tokens. Humaneval is known to cause degeneration after given stopwords, to limit potential faulty generation of new tokens to we set max new tokens to 400. To ensure long enough reasoning chains we limited the generation on StrategyQA to 450 tokens.
Appendix J Abstract consistency
J.1 Temperature sets
@@ I.4 Token generation
J.2 Weighing abstract consistency
@@
K.1 Used k-shot prompts
@@
A: Leah had 32 chocolates and Leah’s sister had 42. That means there were originally 32 + 42 = 74 chocolates. 35 have been eaten. So in total they still have 74 - 35 = 39 chocolates. The answer is 39.
Q: Jason had 20 lollipops. He gave Denny some lollipops. Now Jason has 12 lollipops. How many lollipops did Jason give to Denny?
@@
A: Michael initially had 58 balls. He lost 23 on Tuesday, so after that he has 58 - 23 = 35 balls. On Wednesday, he lost 2 more, so now he has 35 - 2 = 33 balls. The answer is 33.
@@
A: She bought 5 bagels for \$3 each. This means she spent 5 * \$3 = \$15 on the bagels. She had \$23 in the beginning, so now she has \$23 - \$15 = \$8. The answer is 8.
  Proposed 4-shot on AQuA-RAT:
@@
A: If 10 is added to each number, then the mean of the numbers also increases by 10. So the new mean would be 50. The answer is (a).
@@
Q: A person is traveling at 20 km/hr and reached his destination in 2.5 hr. Find the distance.
@@
Proposed 6-shot on StrategyQA:  [^4] For commonsense reasoning, we parse the full string answer as the final answer after the model generates “The answer is ”
Q: Do hamsters provide food for any animals?
A: Hamsters are prey animals. Prey are food for predators. Thus, hamsters provide food for some animals. So the answer is yes.
Q: Could Brooke Shields succeed at University of Pennsylvania?
A: Brooke Shields went to Princeton University. Princeton University is about as academically rigorous as the University of Pennsylvania. Thus, Brooke Shields could also succeed at the University of Pennsylvania. So the answer is yes.
Q: Yes or no: Hydrogen’s atomic number squared exceeds number of Spice Girls?
A: Hydrogen has an atomic number of 1. 1 squared is 1. There are 5 Spice Girls. Thus, Hydrogen’s atomic number squared is less than 5. So the answer is no
Q: Yes or no: Is it common to see frost during some college commencements?
A: College commencement ceremonies can happen in December, May, and June. December is in the winter, so there can be frost. Thus, there could be frost at some commencements. So the answer is yes.
Q: Yes or no: Could a llama birth twice during War in Vietnam (1945-46)?
A: The War in Vietnam was 6 months. The gestation period for a llama is 11 months, which is more than 6 months. Thus, a llama could not give birth twice during the War in Vietnam. So the answer is no.
Q: Yes or no: Would a pear sink in water?
A: The density of a pear is about 0.6g/cm³, which is less than water. Objects less dense than water float. Thus, a pear would float. So the answer is no.
The reluctance to follow prompt formatting, primarily observed in GPT-3.5-turbo and GPT-4o-mini on certain tasks, stems from differences in fine-tuning and intended usage. Specifically, "chat fine-tuning" causes models to disregard formatting guidelines provided by few-shot prompts. To address this issue, we include information to keep the formatting uniform for chat fine-tuned models.
Appendix L Datasets
We selected the datasets that are commonly used in similar methods such as baseline self-consistency [33] and related work to simplify reproduction and comparison to ensure consistency in our results.
We use the recommended configuration splits for testing as suggested by default for each dataset. For AQuA-RAT, our test set includes the full set of 254 examples. In the case of StrategyQA, we employ the complete test split, which consists of 687 samples. Specifically, for SVAMP, we utilize the train and test split comprising 1,000 samples to achieve a less noisy evaluation.
Appendix M K-means Clustering
@@
M.1.2 Average correct datapoint proportion
@@
M.1.3 Cluster density comparison
@@
SVAMP GPT 3.5 4.428 4.513
SVAMP GPT 4o mini 4.356 4.653
SVAMP LLAMA 3 4.562 4.569
AQuA-RAT LLAMA 2 0.838 0.670
@@
AQuA-RAT GPT 4o mini 2.134 3.082
AQuA-RAT LLAMA 3 3.235 3.163
StrategyQA LLAMA 2 2.741 3.215
StrategyQA Mistral 1.962 2.487
StrategyQA GPT 3.5 4.283 4.751
StrategyQA GPT 4o mini 1.869 2.935
StrategyQA LLAMA 3 2.864 3.124
Appendix N Abstract consistency on different temperature sets
@@
Appendix O Dimensionality reduction